# DER Day-Ahead Offering: A Neural Network Column-and-Constraint Generation Approach

### v1

@@ Abstract
In the day-ahead energy market, the offering strategy of distributed energy resource (DER) aggregators must be submitted before the uncertainty realization in the form of price-quantity pairs. This work addresses the day-ahead offering problem through a two-stage robust adaptive stochastic optimization model, wherein the first-stage price-quantity pairs and second-stage operational commitment decisions are made before and after DER uncertainty is realized, respectively. Uncertainty in day-ahead price is addressed using a stochastic programming, while uncertainty of DER generation is handled through robust optimization. To address the max-min structure of the second-stage problem, a neural network-accelerated column-and-constraint generation method is developed. A dedicated neural network is trained to approximate the value function, while optimality is maintained by the design of the network architecture. Numerical studies indicate that the proposed method yields high-quality solutions and is up to $100$ times faster than Gurobi and $33$ times faster than classical column-and-constraint generation on the same 1028-node synthetic distribution network.
@@ I Introduction
With the increasing penetration of distributed energy resources (DERs), DER aggregators face new challenges in in participating in day-ahead electricity markets with significant generation and price uncertainty [iea_unlocking_2022]. An offering strategy defines how an aggregator determines the quantity and price of energy or flexibility to be offered in the market while satisfying the operational and network constraints of the aggregated resources. Significant attention has been drawn to the design of economic day-ahead offering strategies through optimization-based methods, including the use of robust optimization or stochastic programming [kaya_fifty_2026].
@@
In current industry practice, day-ahead DER offering is typically formulated as a two-stage adaptive robust optimization to capture price and generation uncertainty: the aggregator determines the “here-and-no” decisions (offering price-quantity pairs) before price and DER uncertainty realization and optimizes the “wait-and-see” recourse decisions (DER dispatch commands) after the uncertainty is revealed. However, robust optimization can be overly conservative in certain settings [baringo_stochastic_2017]. By contrast, stochastic programming represents uncertainty through an explicit set of sampled scenarios. Its disadvantages lie in the requirement of the knowledge of explicit probabilistic distribution of the uncertain parameters and might lead to the so-called “curse of dimensionality” in practice [plazas_multimarket_2005]. To well balance robust optimization and stochastic programming, the application of two-stage hybrid adaptive robust stochastic optimization (2S-ARSO) becomes attractive for DER day-ahead offering. With the advancement of optimization methods, existing approaches to solve 2S-ARSO include: 1) approximation algorithms [bertsimas_theory_2010]; 2) Benders decomposition [bertsimas_adaptive_2013]; 3) column-and-constraint generation (CCG) [zeng_solving_2013]; and 4) scenario reduction method [sun_robust_2021]. However, these methods may not scale well for large-scale problems. Given the computational burden stemming from the max-min recourse structure and the large scenario sets inherent to 2S-ARSO, its practical deployment is often constrained, particularly in industrial environments where operational time and computational resources are limited.
@@
In recent years, growing interest has focused on embedding neural network (NN) models within classical algorithmic pipelines to improve efficiency and scalability in practice [zhang_survey_2023]. Much of the machine learning (ML) for linear program (LP) literature targets deterministic settings, so adopting these techniques to 2S-ARSO is nontrivial. The core difficulty is its intrinsic infinite-dimensional characteristic: the recourse (wait-and-see) decisions are decision rules defined over the uncertainty space and parameterized by the here-and-now decisions, yielding a functional dependence that enlarges the feasible set beyond finite-dimensional formulations. Second, 2S-ARSO embeds a special tri-level structure: choose the here-and-now decisions, identify the worst-case uncertainty conditional on them, and optimize the ensuing wait-and-see recourse. Any practical method must address all three components in a unified manner. Third, the computational burden of solving the 2S-ARSO is negligible when considering network constraints and a large number of uncertainty scenarios. Finally, during the iteration of CCG, the induced optimization becomes progressively harder and may become computationally intractable as iteration number grows.
@@
To address these challenges, we develop a tailored NN-accelerated CCG framework that relieves the computational burden of 2S-ARSO for DER day-ahead offering strategy design. The NN-based CCG algorithm is based on [dumouchelle_neur2ro_2024], while we adapt it to our specific problem setting. Specifically, a NN is trained to estimate the optimal cost of the value function of the recourse problem and to identify the corresponding worst-case uncertainty realization. The resulting predictions tighten the large-scale, effectively infinite-dimensional linear recourse and streamline the classical CCG procedure (in Section III-B). To the best of our knowledge, this is the first study leveraging ML to address the 2S-ARSO-based DER day-ahead offering problem. The training period is implemented once, and the trained networks can be easily utilized for day-ahead offering decision making while achieving significant speed-ups that outperform the state-of-the-art commercial solvers. Our key contributions are summarized as follows:
@@
    We design a tailored neural architecture that jointly estimates worst-case realizations and the recourse value, which is invariant to problem dimension.
@@
    We demonstrate up to $100$-fold speedup for the NN-accelerated CCG relative to conventional CCG on synthetic networks with real historical price data.
The rest of the paper is organized as follows. Section II introduces the compact formulation of the day-ahead offering problem as a 2S-ARSO. Section III proposes the NN-accelerated CCG algorithm for solving the 2S-ARSO-based day-ahead offering problem. Numerical studies are included in Section IV. Section V concludes the paper.
@@ II Hybrid Adaptive Robust Stochastic Optimization Formulation
In this work, we consider the day-ahead offering problem for a DER aggregator over 24 hours. The aggregator submits price-quantity offers (first-stage, “here-and-now”) before DER uncertainty is realized. After DER uncertainties are revealed, the aggregator takes operational recourse (second-stage, “wait-and-see”) decisions that includes DER dispatch, battery operations, and power flow constraints.
We model the 2S-ARSO day-ahead offering problem as a two-stage adaptive robust optimization (2S-ARO) problem under sampled price trajectories. For the 2S-ARO model, the decision vector is split into “here-and-now” variables ($\boldsymbol{x}$) and “wait-and-see” (recourse) variables ($\boldsymbol{y}$). The here-and-now decisions (submitted price-quantity pairs) are fixed before DER uncertainty is realized, whereas the cost-minimizing recourse decisions (DER dispatch and other real-time decisions) are determined after the uncertainty is revealed. The day-ahead offering model is formulated as the following risk-neutral expectation over all sampled day-ahead price scenarios:
@@
In the problem above, the day-ahead price at the first stage is modeled by stochastic price trajectories $\boldsymbol{\omega}\in\Omega$ sampled through the Markov process (each trajectory is assigned a nonnegative weight $\rho_{\omega}$) [zheng_arbitraging_2022]. The DER uncertainty is modeled through the budgeted uncertainty set $\boldsymbol{\xi}\in\mathcal{U}$ [thiele_robust_2010]. The first-stage cost $\boldsymbol{c}^{\top}(\boldsymbol{\omega})\boldsymbol{x}$ describes the negative value of the day-ahead offering revenue given the price uncertainty $\boldsymbol{\omega}$. The recourse cost $\boldsymbol{b}^{\top}\boldsymbol{y}$ represents the cost of real-time DER operation given the realization of DER uncertainty $\boldsymbol{\xi}$. The feasible regions of the first- and second-stage problems are denoted by $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$, respectively. To be consistent with U.S. electricity market regulations [kim_benefits_2021], our framework discourages inter-settlement arbitrage between two-settlement markets. This is done by modeling the real-time price as a small deviation from the day-ahead one in such a way that penalizes any intentional real-time power deviation.
The feasible region $\mathcal{X}$ of the first-stage problem is a bounded polyhedron (polytope) given by:
@@
where (2) contains both the monotonicity constraint of offering curves and the capacity constraints of offering quantities. The feasible region $\mathcal{Y}(\boldsymbol{x},\boldsymbol{\xi})$ of the second-stage problem can be characterized by:
@@
where (3) includes DER power availability constraint, physical constraint of DERs, and linear power flow constraints [gan_convex_2014].
@@ III Neural Network-Accelerated Column-and-Constraint Generation Algorithm
The problem in (5) identifies the worst-case scenario $\boldsymbol{\xi}$ over the inner minimization problem. Given the polyhedral structure of $\mathcal{Y}$ and the right-hand side (RHS) uncertainty of $\boldsymbol{\xi}$, under the relatively complete recourse assumption, the inner minimization problem is a convex piecewise linear function in $\boldsymbol{\xi}$, whose maximization over a polytopic uncertainty set $\mathcal{U}$ is attained at one of its extreme points [sun_robust_2021, Sect. 3.2.2].
@@ III-A Column-and-Constraint Generation Algorithm
The CCG algorithm solves problem (4) by exploiting the structural insights mentioned above in an iterative master problem/subproblem fashion. A detailed description of the algorithm can be found in [sun_robust_2021, Sect. 3.2.6]. We provide a very brief introduction here to facilitate further development.
The CCG algorithm can be viewed as a constraint generation algorithm: for a given first-stage candidate solution $\boldsymbol{x}^{(k)}$, it identifies a worst-case scenario $\boldsymbol{\xi}^{(k)}\in\operatorname{ext}(\mathcal{U})$ for the second-stage problem and generates a set of new constraints for the master problem, where $\operatorname{ext}(\cdot)$ denote the set of extreme points of a polytope. At iteration $k$, the master problem of (4) is given by:
$\min\quad$ $\boldsymbol{\eta}$ (6a)
@@
s.t. $\boldsymbol{\eta}\;\geq\;\boldsymbol{c}(\boldsymbol{\omega})^{\top}\boldsymbol{x}\;+\;\boldsymbol{b}^{\top}\boldsymbol{y}^{(i)},$ $i=1,\ldots,k$ (6b)
@@
where $\boldsymbol{\xi}^{(i)},i=1,\ldots,k$ are the extreme points of $\mathcal{U}$. It is straightforward to see that problem (6) is a relaxation of (4) and thus provides a lower bound to (4).
To certify the validity of the current first-stage solution $\boldsymbol{x}^{(k)}$ and to identify the worst-case violation, we formulate a subproblem that computes the worst-case realization of the uncertainty for $\boldsymbol{x}^{(k)}$. By invoking the strong duality of LP, the inner minimization in (5) admits the following equivalent reformulation:
@@
where $\boldsymbol{\pi}^{(1)},\ldots,\boldsymbol{\pi}^{(r)}$ and $\boldsymbol{\xi}^{(1)},\ldots,\boldsymbol{\xi}^{(m)}$ are the extreme points of $\mathcal{P}:=\{\boldsymbol{\pi}\geq 0:\boldsymbol{F}^{\top}\boldsymbol{\pi}=\boldsymbol{b}\}$ and $\mathcal{U}$, respectively. Since both feasible sets $\mathcal{P}$ and $\mathcal{U}$ are convex polytopes, the worst-case value $\mathcal{Q}(\boldsymbol{x})$ is attained at their extreme points, which allows the subproblem (7) to be represented as a finite maximization over the extreme points of $\mathcal{P}$ and $\mathcal{U}$. Given the current candidate first-stage solution $\boldsymbol{x}^{(k)}$, the corresponding upper bound $\mathrm{UB}$ to (4) is updated as
@@ III-B NN-Accelerated Column-and-Constraint Generation
Although $\mathcal{Q}\left(\boldsymbol{x}\right)$ is a convex piecewise linear function of $\boldsymbol{x}$, evaluating its value for any fixed $\boldsymbol{x}$ is a hard problem due to the presence of bilinear terms in (7) which destroys convexity in $\boldsymbol{\pi}$ and $\boldsymbol{\xi}$. The computational burden of (7) becomes even worse with the consideration of power flow constraints, which is normally ignored [baringo_stochastic_2017] but considered herein. We develop a neural estimator that, given a first-stage decision, predicts the optimal objective value of the corresponding second-stage problem with specific uncertainty realization with high accuracy. The NN is then integrated within CCG to accelerate the solution of both the master problem and the subproblem. It does so using one mixed-integer linear program (MILP)-representable NN with learnable parameters $\Theta$ that can assess, for a first-stage decision $\boldsymbol{x}$ under uncertain $\boldsymbol{\xi}$, the following:
@@ III-B2 Subproblem
Since $\mathcal{U}$ is polytopic, (11) is an MILP which is solvable by off-the-shelf solvers.
To determine whether a new $\boldsymbol{\xi}^{*}\in\mathcal{U}$ has been identified, we compare its prediction to those already used in the master problem:
@@
As the iteration count rises, evaluations of the neural modules impose increasing computational overhead. We therefore adopt an architecture that can implement NN-accelerated CCG both efficiently and effectively. After sampling historical data, we embed the first-stage decisions $(\boldsymbol{x})$ and DER uncertainty realizations $(\boldsymbol{\xi})$ using dedicated low-dimensional embedding NNs $(\Phi_{\boldsymbol{x}}\text{ and }\Phi_{\boldsymbol{\xi}})$ for each instance. The resulting embeddings are concatenated and fed to the value network $\Phi_{\Theta^{\ast}}$, which estimates the worst-case scenario and the optimal recourse objective, and serves as an oracle module within the pipeline for subsequent optimization. The neural architecture for solving the recourse problem is illustrated in Figure 1. For the master problem, since the scenario parameters are not variables here, the scenario embeddings $\Phi_{\xi}(\xi^{(k)})$ can be precomputed via a forward pass for each scenario. In the finalized master problem, only the decision-embedding network $\Phi_{\boldsymbol{x}}$ and the value NN $\Phi_{\Theta^{*}}$ are required. Both can be implemented with standard ReLU-activated networks that lead to MILP formulations. In addition, duplicating these modules across scenarios remains computationally tractable. For the second-stage recourse problem, it suffices to encode only the decision embedding $\Phi_{\boldsymbol{x}}$ together with the value network $\Phi_{\Theta^{\ast}}$, since $\Phi_{\boldsymbol{x}}(\boldsymbol{x})$ can be precomputed efficiently via a single forward pass.
To generalize across instances, the architecture should be invariant to variable cardinality and ordering and robust to permutations of constraint/objective coefficients. We can enforce these properties with set-based NNs for $\Phi_{\boldsymbol{x}}$ and $\Phi_{\Theta^{\ast}}$, which yield permutation-invariant encoders [zaheer_deep_2018]. Detailed generalization process includes: Before feeding first-stage solution and DER uncertainty scenarios to the embedding networks, these variables are decomposed into per-variable representations containing their values, constraint coefficients, and objective weights. The resulting set of per-variable embeddings is then aggregated (e.g., sum pooling) and passed through a feedforward network to produce the first-stage and scenario representations. Since the supplemented architecture of set-based NNs is not the key innovative component of this work, we omit further architectural details and refer to this component as the “Encoder” in Fig. 1, where the corresponding arrow is highlighted in red. The detailed description of the NN-accelerated CCG algorithm is summarized in Algorithm 1.

### v2

@@ Abstract
In the day-ahead energy market, the offering strategy of distributed energy resource (DER) aggregators must be submitted before the uncertainty realization in the form of price-quantity pairs. This work addresses the day-ahead offering problem through a two-stage adaptive robust stochastic optimization model, wherein the first-stage price-quantity pairs and second-stage operational commitment decisions are made before and after DER uncertainty is realized, respectively. Uncertainty in day-ahead price is addressed using a stochastic programming-based approach, while uncertainty of DER generation is handled through robust optimization. To address the max-min structure of the second-stage problem, a neural network-accelerated column-and-constraint generation method is developed. A dedicated neural network is trained to approximate the value function, while optimality is maintained by the design of the network architecture. Numerical studies indicate that the proposed method yields high-quality solutions and is up to $100$ times faster than Gurobi and $33$ times faster than classical column-and-constraint generation on the same 1028-node synthetic distribution network.
@@ I Introduction
With the increasing penetration of distributed energy resources (DERs), DER aggregators face new challenges in participating in day-ahead electricity markets with significant generation and price uncertainty [7]. An offering strategy defines how an aggregator determines the quantity and price of energy or flexibility to be offered in the market while satisfying the operational and network constraints. Significant attention has been drawn to the design of economic day-ahead offering strategies through optimization-based methods, including the use of robust optimization or stochastic programming [8].
@@
Currently, the day-ahead DER offering problem is typically formulated as a two-stage adaptive robust optimization to capture price and generation uncertainty: the aggregator determines the “here-and-now” decisions (offering price-quantity pairs) before the realization of price and DER uncertainty and optimizes the “wait-and-see” recourse decisions (DER dispatch commands) after the uncertainty is revealed. However, robust optimization can be overly conservative in certain settings [1]. By contrast, stochastic programming-based approaches model uncertainty through an explicit set of sampled scenarios. Its disadvantages lie in the requirement for knowledge of the explicit probabilistic distribution of the uncertain parameters, which might lead to the so-called “curse of dimensionality” [10]. To balance both approaches effectively, the application of two-stage adaptive robust stochastic optimization (2S-ARSO) becomes attractive. With the advancement of optimization methods, existing approaches to solve 2S-ARSO include: 1) approximation algorithms [2]; 2) Benders decomposition [3]; 3) column-and-constraint generation (CCG) [14]; and 4) scenario reduction method [11]. However, these methods may not scale well for large-scale problems. Given the computational burden stemming from the max-min recourse structure and the large scenario sets inherent to 2S-ARSO, its practical deployment is often constrained, particularly in industrial environments where operational time and computational resources are limited.
@@
In recent years, growing interest has focused on embedding neural network (NN) models within classical algorithmic pipelines to improve efficiency and scalability in practice [15]. Much of the machine learning (ML) literature for linear programming (LP) targets deterministic settings, making it inapplicable to our 2S-ARSO problem. The core difficulty lies in its intrinsic infinite-dimensional characteristic: the recourse decisions are decision rules defined over the uncertainty space and parameterized by the here-and-now decisions, yielding an infinite-dimensional functional dependence. Second, 2S-ARSO embeds a special tri-level structure: choosing a here-and-now decision, identifying the worst-case uncertainty realization, and optimizing the ensuing recourse decision. Any practical method must address all three components in a unified manner. Third, the computational burden of solving the 2S-ARSO is negligible when considering network constraints and a large number of uncertainty scenarios. Finally, during the iteration of CCG, the induced optimization becomes progressively harder and may become computationally intractable as the number of iterations grows.
@@
To address these challenges, we develop a tailored NN-accelerated CCG framework that relieves the computational burden of 2S-ARSO for DER day-ahead offering strategy design. The NN-based CCG algorithm is based on [4], while we adapt it to our specific problem setting. Specifically, an NN is trained to estimate the optimal cost of the value function of the recourse problem and to identify the corresponding worst-case uncertainty realization. The resulting predictions eliminate the need to solve the infinite-dimensional linear recourse problem and streamline the classical CCG procedure (see Section III-B). To the best of our knowledge, this is the first study leveraging ML to address the 2S-ARSO-based DER day-ahead offering problem. The training process is conducted once, and the trained NNs can be easily utilized for day-ahead offering decision making while achieving significant speed-ups that outperform the state-of-the-art commercial solvers. Our key contributions are summarized as follows:
@@
    We design a tailored problem dimension-invariant neural architecture that jointly estimates worst-case realizations and the recourse value.
@@
    We demonstrate up to $100$-fold speedup of the NN-accelerated CCG over conventional CCG on synthetic networks using historical price data.
The rest of the paper is organized as follows. Section II introduces the compact formulation of the day-ahead offering problem as a 2S-ARSO. Section III proposes the NN-accelerated CCG solution algorithm. Numerical studies are included in Section IV. Section V concludes the paper.
@@ II Hybrid Adaptive Robust Stochastic Optimization Formulation
In this work, we consider the day-ahead offering problem for a DER aggregator over 24 hours. The aggregator submits price-quantity offers (first-stage, “here-and-now”) before DER uncertainty is realized. After DER uncertainties are revealed, the aggregator makes operational recourse decisions (second-stage, “wait-and-see”) regarding DER dispatch, which are constrained by device and power flow constraints.
We model the 2S-ARSO day-ahead offering problem as a 2S-ARO problem under sampled price trajectories. For the 2S-ARO model, the decision vector is split into “here-and-now” variables ($\boldsymbol{x}$) and “wait-and-see” (recourse) variables ($\boldsymbol{y}$). The here-and-now decisions (submitted price-quantity pairs) are fixed before the DER uncertainty is realized, whereas the cost-minimizing recourse decisions (DER dispatch and other real-time decisions) are determined after the uncertainty is revealed. The day-ahead offering model is formulated as the following risk-neutral expectation over all sampled day-ahead price scenarios:
@@
In the problem above, the day-ahead price at the first stage is modeled by stochastic price trajectories $\boldsymbol{\omega}\in\Omega$ sampled through the Markov process (each trajectory is assigned a nonnegative weight $\rho_{\omega}$) [16]. The DER uncertainty is modeled through the budgeted uncertainty set $\boldsymbol{\xi}\in\mathcal{U}$ [12]. The first-stage cost $\boldsymbol{c}^{\top}(\boldsymbol{\omega})\boldsymbol{x}$ describes the negative value of the day-ahead offering revenue given the price uncertainty $\boldsymbol{\omega}$. The recourse cost $\boldsymbol{b}^{\top}\boldsymbol{y}$ represents the cost of real-time DER operation given the realization of DER uncertainty $\boldsymbol{\xi}$. The feasible regions of the first- and second-stage problems are denoted by $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$, respectively. To be consistent with U.S. electricity market regulations [9], our framework discourages inter-settlement arbitrage between two-settlement markets. This is done by modeling the real-time price as a small deviation from the day-ahead one in such a way that penalizes any intentional real-time power deviation.
The feasible region of the first-stage problem $\mathcal{X}$ is a bounded polyhedron (polytope) given by
@@
which contains both the monotonicity constraint of offering curves and the capacity constraints of offering quantities. The feasible region of the second-stage problem $\mathcal{Y}(\boldsymbol{x},\boldsymbol{\xi})$ is
@@
which includes DERs power availability constraints, DER physical constraints, and linear power flow constraints [5].
@@ III Neural Network-Accelerated Column-and-Constraint Generation Algorithm
The problem in (5) identifies the worst-case scenario $\boldsymbol{\xi}$ over the inner minimization problem. Given the polyhedral structure of $\mathcal{Y}$ and the right-hand side (RHS) uncertainty of $\boldsymbol{\xi}$, under the relatively complete recourse assumption, the inner minimization problem is a convex piecewise linear function in $\boldsymbol{\xi}$, whose maximization over a polytopic uncertainty set $\mathcal{U}$ is attained at one of its extreme points [11, Sect. 3.2.2].
@@ III-A Column-and-Constraint Generation Algorithm
The CCG algorithm solves problem (4) by exploiting the structural insights mentioned above in an iterative master problem/subproblem fashion. A detailed description of the algorithm can be found in [11, Sect. 3.2.6]. We provide a very brief introduction here to facilitate further development.
The CCG algorithm can be viewed as a constraint generation algorithm: for a given first-stage candidate solution $\boldsymbol{x}^{(k)}$, it identifies a worst-case scenario $\boldsymbol{\xi}^{(k)}\in\operatorname{ext}(\mathcal{U})$ for the second-stage problem and generates a set of new constraints for the master problem, where $\operatorname{ext}(\cdot)$ denote the set of extreme points of a polytope. At iteration $k$, the master problem of (4) is
$\min\quad$ $\eta$ (6a)
@@
s.t. $\eta\;\geq\;\boldsymbol{c}(\boldsymbol{\omega})^{\top}\boldsymbol{x}\;+\;\boldsymbol{b}^{\top}\boldsymbol{y}^{(i)},$ $i=1,\ldots,k$ (6b)
@@
where $\boldsymbol{\xi}^{(i)}\in\operatorname{ext}(\mathcal{U})$. It is straightforward to see that problem (6) is a relaxation of (4) and thus provides a lower bound.
To certify the validity of the current first-stage solution $\boldsymbol{x}^{(k)}$ and to identify the worst-case violation, we formulate a subproblem that computes the worst-case uncertainty realization for $\boldsymbol{x}^{(k)}$. By invoking LP strong duality, the inner minimization in (5) admits the following equivalent reformulation:
@@
where $\boldsymbol{\pi}^{(j)}$ are the extreme points of $\mathcal{P}:=\{\boldsymbol{\pi}\geq 0:\boldsymbol{F}^{\top}\boldsymbol{\pi}=\boldsymbol{b}\}$. Since both $\mathcal{P}$ and $\mathcal{U}$ are convex polytopes, the worst-case value $\mathcal{Q}(\boldsymbol{x})$ is attained at one of their extreme points, which allows (7) to be represented as a finite maximization over $\operatorname{ext}(\mathcal{P})$ and $\operatorname{ext}(\mathcal{U})$. Given the current first-stage solution $\boldsymbol{x}^{(k)}$, the corresponding upper bound $\mathrm{UB}$ to (4) is updated as
@@ III-B NN-Accelerated Column-and-Constraint Generation
Although $\mathcal{Q}\left(\boldsymbol{x}\right)$ is a convex piecewise linear function of $\boldsymbol{x}$, evaluating its value for any fixed $\boldsymbol{x}$ is a hard problem due to the presence of bilinear terms in (7) which destroys convexity in $\boldsymbol{\pi}$ and $\boldsymbol{\xi}$. The computational burden of (7) becomes even worse with the consideration of power flow constraints, which is normally ignored [1] but considered herein. We develop a neural estimator that, given a first-stage decision, predicts the optimal objective value of the corresponding second-stage problem with specific uncertainty realization with high accuracy. The NN is then integrated within CCG to accelerate the solution of both the master problem and the subproblem. It does so using one mixed-integer linear program (MILP)-representable NN with learnable parameters $\Theta$ that can assess, for a first-stage decision $\boldsymbol{x}$ under uncertain $\boldsymbol{\xi}$, the following:
@@ III-B2 Subproblem
Since $\mathcal{U}$ is polytopic, (11) is an MILP that is solvable by off-the-shelf solvers.
To determine whether a new $\boldsymbol{\xi}^{*}\in\mathcal{U}$ has been identified, we compare its prediction to those already in the master problem:
@@
As the iteration count rises, evaluations of the neural modules impose increasing computational overhead. We therefore adopt an architecture that can implement NN-accelerated CCG both efficiently and effectively. After sampling historical data, we embed $\boldsymbol{x}$ and $\boldsymbol{\xi}$ using dedicated low-dimensional embedding NNs $(\Phi_{\boldsymbol{x}}\text{ and }\Phi_{\boldsymbol{\xi}})$ for each instance. The resulting embeddings are concatenated and fed to the value network $\Phi_{\Theta^{\ast}}$, which estimates the worst-case scenario and the optimal recourse objective and serves as an oracle module within the pipeline for subsequent optimization. The neural architecture for solving the recourse problem is shown in Figure 1. For the master problem, since the scenario parameters are fixed, the scenario embeddings $\Phi_{\xi}(\xi^{(k)})$ can be precomputed via a forward pass for each scenario. In the finalized master problem, only the decision-embedding network $\Phi_{\boldsymbol{x}}$ and the value NN $\Phi_{\Theta^{*}}$ are required. Both can be implemented with standard ReLU-activated networks that lead to MILP formulations. In addition, duplicating these modules across scenarios remains computationally tractable. For the second-stage recourse problem, it suffices to encode only the decision embedding $\Phi_{\boldsymbol{x}}$ together with the value network $\Phi_{\Theta^{\ast}}$, since $\Phi_{\boldsymbol{x}}(\boldsymbol{x})$ can be precomputed efficiently via a single forward pass.
@@
To generalize across instances, the architecture should be invariant to variable cardinality and ordering; and robust to permutations of constraint/objective coefficients. We enforce these properties by set-based NNs with shared parameters, which yield permutation-invariant encoders [13] and improve generalization to unseen instances. The detailed generalization process includes the following: Before feeding the first-stage solution and DER uncertainty scenarios to the embedding networks, these variables are decomposed into per-variable representations containing their values, constraint coefficients, and objective weights. To avoid underestimating the worst-case recourse cost, we check each NN-identified candidate scenario by solving the exact second-stage recourse problem. If it violates the tolerance, we add it to the active set and continue. The resulting set of per-variable embeddings is then aggregated (e.g., sum pooling) and passed through a feedforward network to produce the first-stage and scenario representations. Due to space limitations, we omit further architectural details and refer to this component as the “Encoder” in Fig. 1, where the corresponding arrow is highlighted in red. The detailed description of the NN-accelerated CCG algorithm is summarized in Algorithm 1. The finite-termination property of the algorithm is established in the Appendix.